OpenStudy (anonymous):

Find the roots of the polynomial equation. x^2 + x - 72 = 0 A. 9 and –8 B. –6 and 6 C. –2 and 16 D. –9 and 8

OpenStudy (hyper):

As this is a quadratic function, it could be solved by using the quadratic formula or through factorisation. As factorisation is simpler to type, x^2 + x - 72 = (x - 8) (x + 9) =0. From here, let x - 8 = 0 and x + 9 = 0. Hence, x = 8 and x = -9. This answer is confirmed by resubbing these values into the quadratic to verify. Hence, the answer is D